Grade 11Mechanics

a truck travelling along a straight road at a constant speed of 72 kmph passes a car at time t=0 which is moving much slower . at the instant the truck passes the car the car starts accelerating at constant 1m/s and over takes the truck 0.6km further down the road the road , form where the car moves uniformly . find the distance between them at time t=50s

YASH
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
For truck,
when car overtakes 
at constannt speed:
t1 = x/72 …..........(1)
 
For car,
x = 0 + 0.5*1*t2^2
t2 = root(2x) …...........(2)
 
x = 0.6 km (given)
t1 = 0.6/72  = 1/120 hr.
t2 = root(1.2) hr.
 
at t= 50s
x1 = (50/60*60)*72
x2 = (50/60*60)^2/2
 
Distance b/w them: 
 |x2- x1|